OpenStudy (angeleyes108):

HELP!! MULTIPLE CHOICE: FAN & MEDAL-What is the circumference of the given circle in terms of pi?

OpenStudy (angeleyes108):



OpenStudy (angeleyes108):

My choices are: 14\[14\pi / 28\pi / 42\pi / 196\pi \]

OpenStudy (anonymous):

Do you know the equation of circle's circumference ? It's 2πr where r is the radius of the circle Let me know if you got it :)

OpenStudy (angeleyes108):

Doing that equation I don't get the answer it's looking for. I listed the choices above. It's in terms of pi.

OpenStudy (anonymous):

What is you given raduis ?

OpenStudy (angeleyes108):

14 in.

OpenStudy (anonymous):

Then plugging it into equation what you will get ?

OpenStudy (angeleyes108):

87.92 and that's not one of my choices.

OpenStudy (anonymous):

No you want it in terms of pi then you don't type the π onto your calculator :)

OpenStudy (anonymous):

You got it ? :)

OpenStudy (angeleyes108):

I'm confused....

OpenStudy (angeleyes108):

that doesn't make sense...

OpenStudy (johnweldon1993):

So the equation is \[\large C = 2 \pi r\] So we just need to deal with the 'r' and the 2 The radius given is 14 inches...so we have \[\large C = 2 \pi \times 14\] or \[\large 28 \pi\] It is in terms of pi

OpenStudy (angeleyes108):

Oh thank you so much!!